### v2

# CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation

Abstract
With the rapid advancement of Large Language Models (LLMs), the demand for robust instruction-following capabilities in code generation tasks has grown significantly. Code generation not only facilitates faster prototyping and automated testing, but also augments developer efficiency through improved maintainability and reusability of code. In this paper, we introduce CodeIF, the first benchmark specifically designed to assess the abilities of LLMs to adhere to task-oriented instructions within diverse code generation scenarios. CodeIF encompasses a broad range of tasks, including function synthesis, error debugging, algorithmic refactoring, and code explanation, thereby providing a comprehensive suite to evaluate model performance across varying complexity levels and programming domains. We conduct extensive experiments with LLMs, analyzing their strengths and limitations in meeting the demands of these tasks. The experimental results offer valuable insights into how well current models align with human instructions, as well as the extent to which they can generate consistent, maintainable, and contextually relevant code. Our findings not only underscore the critical role that instruction-following LLMs can play in modern software development, but also illuminate pathways for future research aimed at enhancing their adaptability, reliability, and overall effectiveness in automated code generation.  [^1] CodeIF data and code are publicly available: https://github.com/lin-rany/codeIF.
Machine Learning, ICML
1Beihang University, 2Tsinghua University, 3National University of Singapore
footnotetext: *Equal contributionfootnotetext: Corresponding Author
1 Introduction
With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation. While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge (Liu et al., 2024; Zhang et al., 2023). To drive progress in this field, a comprehensive and systematic evaluation framework is essential (Jiang et al., 2024; Zhou et al., 2023).
This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation. Built upon insights from existing evaluation sets like McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++. It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
To ensure rigorous assessment, we propose four novel evaluation metrics: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics measure models’ ability to handle multi-constraint problems from different perspectives, including full compliance, average constraint satisfaction, logical coherence, and consistency in instruction execution. By offering a structured evaluation framework, CodeIF provides valuable insights into the current state and future direction of LLM-driven code generation.
Overall, our contributions are mainly four-fold:
  1. 1.
    Innovative Benchmark: We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation. CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
  2. 2.
    Automated High-Quality Instruction Generation: Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists. This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
  3. 3.
    Novel Evaluation Metrics: We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks. These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
  4. 4.
    Extensive Evaluation and Analysis: We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels. Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
2 CodeIF
Overview:
Figure 1: The construction process of CodeIF. The first step involves the construction of constraint instructions, followed by the assembly of the dataset, and finally the construction of dependencies between instructions.
As shown in Figure 1, the construction of CodeIF includes the following steps: we first collect and refine constraint instructions from real code generation tasks, then we combine code generation tasks with LLMs and human review to create a high-quality evaluation dataset.
2.1 Building
The construction of the CodeIF dataset consists of two main phases: collection of constraint instructions and data processing. In Section  2.2, we elaborate on how we collected the constraint instructions for code generation. Following this, in Section  2.3, we describe how we combined these constraint instructions to produce the final CodeIF evaluation dataset.
2.2 Constraint Instructions Collection
The first phase of our work focuses on code generation by collecting and formulating constraint instructions to construct the CodeIF evaluation dataset. This process consists of two stages: (1) collecting and verifying constraint instructions and (2) applying them to dataset generation.
We begin with an in-depth analysis of code generation tasks based on existing benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), establishing an instruction system covering eight categories. Each category targets a specific aspect of code generation, enabling a fine-grained assessment of LLMs’ instruction-following abilities.
Within each category, we further refine constraints into atomic instructions with clear, explicit directives. These atomic constraints allow binary evaluation (“yes” or “no”), ensuring an objective and precise assessment of model compliance while minimizing subjectivity.
Our classification of atomic constraint instructions is guided by an analysis of existing code generation benchmarks and widely used code repositories. We observe that requirement distribution varies significantly across different levels—line-level, function-level, and full-solution generation. To ensure comprehensive coverage, we design atomic instruction categories that assess both high-level architectural constraints and fine-grained variable control.
To systematically evaluate LLMs’ instruction-following capabilities in code generation, we define eight constraint categories. The Global category ensures that generated code aligns with overarching specifications and holistic constraints. Structural Control focuses on adherence to control structures such as loops and conditionals, as well as data structures like sets and maps. Variable constraints assess details at the variable level, including naming conventions and initialization methods.
At different abstraction levels, Interface, Function, and Class constraints evaluate the model’s ability to construct well-defined program components. The File category tests handling of cross-file dependencies and external library calls. Finally, the Combination category presents a comprehensive challenge by integrating constraints from multiple dimensions, assessing how well models follow complex, multi-faceted instructions in realistic scenarios.
Figure 2: CodeIF Constraints Instruction Distribution
As shown in Figure 2, CodeIF’s distribution across different programming languages and categories is presented.
Through a detailed analysis of existing code generation benchmarks, we designed an evaluation system comprising 8 main categories and 50 fine-grained constraint instructions. This system not only enables a systematic assessment of LLMs’ performance in code generation but also establishes clear benchmarks for future research.
These instructions cover constraints ranging from high-level structural guidelines to fine-grained variable operations, ensuring a comprehensive evaluation. For instance, the Global category assesses a model’s ability to maintain overall structural integrity and logical coherence, while the Variable category evaluates its precision in variable naming and initialization.
By refining constraints into well-defined categories, we can precisely identify the strengths and weaknesses of different LLMs, guiding optimization efforts and advancing the field toward more effective automated code generation. Additionally, this structured evaluation approach provides a valuable reference for future research, fostering continuous innovation. The complete list of atomic constraint instructions is provided in Appendix Table 4.
2.3 Data Construction
Multi-Language and Difficulty-Differentiated Benchmark Design
To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024). These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
To further refine the evaluation, we categorize tasks into two difficulty levels: Easy and Hard. The Hard set includes longer, more intricate instruction lists, designed to rigorously test LLMs’ ability to handle complex constraints.
Automated Generation of Constraint Instructions
To construct high-quality constraint instructions, we leveraged large language models (LLMs) such as GPT-4 to automatically generate task-specific instruction lists. Initially, we curated approximately 20 detailed examples, each featuring a specific code generation task. Human experts then formulated atomic constraint instructions to ensure precise adherence during code generation.
These examples served as prompts, allowing LLMs to generate corresponding instruction lists while refining task descriptions by identifying and eliminating redundant constraints. This approach streamlined instruction-following requirements, ensuring clarity and promoting high-quality outputs.
Constructing Instruction Dependencies
Since some atomic constraints exhibit interdependencies, we further utilized LLMs to construct explicit dependency relationships within instruction sets. These dependencies were integrated into our evaluation framework, enhancing assessment depth and improving the accuracy and practicality of constraint verification.
Specifically, by analyzing the logical connections and sequential dependencies among atomic instructions, we can precisely define the execution steps required for each task. For instance, in function generation tasks, the process typically involves defining the function name first, specifying parameter types, and finally implementing the function body. These dependencies ensure coherence and consistency throughout the code generation process.
Integrating these dependencies into the evaluation system enables a more accurate assessment of a model’s ability to follow complex instructions while providing a foundation for future optimizations. Additionally, this approach helps identify potential challenges models face when handling specific instruction sequences, offering insights for further improvement. As shown in Figure 3, this example from CodeIF illustrates a specific generation task, its associated instruction list, and the corresponding instruction dependencies.
Figure 3: Specific cases of the CodeIF dataset, ’Task’ denotes the specific generation task, ’Type’ refers to the type of constraint, and ’Dependence’ indicates the prerequisite constraints for this constraint.
2.4 Data Analysis
Table 1: CodeIF dataset statistics, showing the statistical information of different difficulty classifications. Avg.Instr represents the average length of the atomic constraint instruction list.
Set Num Avg.Instr Go Python Java C++
Easy 600 11.99 127 165 176 132
Hard 600 13.80 103 183 177 137
Full 1200 12.90 230 348 353 269
CodeIF Static Analysis
Figure 4: The distribution of atomic instruction list lengths in datasets of different difficulties.
As shown in Table 1, the dataset is categorized into three difficulty levels: Easy, Hard, and an aggregated set, Full. The Easy and Hard sets each contain 600 code generation tasks, while the Full dataset combines both, totaling 1,200 tasks. These tasks span four programming languages: Go, Python, Java, and C++.
Among them, Java has the highest number of tasks (353), followed by Python (348). C++ and Go contain 269 and 230 tasks, respectively. The dataset also records the average number of instructions per task for each difficulty level. The Easy set averages 11.99 instructions per task, while the Hard set increases to 13.8, reflecting greater complexity and stricter constraints. The Full dataset, combining both, has an average instruction length of 12.9. Figure 4 illustrates the distribution of task lengths across the dataset.
Table 2: CodeIF dataset statistics information, showing the distribution of atomic restriction instruction categories under different difficulty classifications.
Set Global Structural Control Variable Interface Function Class File Combination
Easy 1638 1008 1336 427 569 544 723 953
Hard 1890 1193 1479 505 659 623 802 1142
Full 3528 2201 2815 932 1228 1167 1525 2095
Figure 5: The distribution of constraint instruction list lengths in datasets of different difficulties.
Analysis of Constraint Instruction Category Distribution
Table 2 presents the distribution of atomic instructions across different difficulty levels in the evaluation set. As shown in Figure 5, we compare the average number of instructions per category between the “Easy” and “Hard” sets. These categories include Global, Structural Control, Variable, Interface, Function, Class, File, and Combination.
Across all categories, the “Hard” dataset contains a higher average number of instructions than the “Easy” set. Notably, the Global category has the highest instruction count in both difficulty levels, averaging 2.5 instructions in the “Easy” set and exceeding 3 in the “Hard” set. This trend indicates that as task complexity increases, models face greater challenges in handling intricate constraints and detailed requirements. By analyzing instruction distribution across categories, we gain deeper insights into model performance under varying levels of difficulty, providing a foundation for further optimization.
Figure 6: Distribution of atomic instruction list lengths across difficulty levels.
Figure 6 illustrates the proportion of each atomic instruction category. The Global category holds the largest share at 22.77%, followed by the Variable category at 18.17%. This distribution highlights CodeIF’s focus on evaluating LLMs’ capabilities at both the overall structural level (e.g., Global constraints) and fine-grained variable handling (e.g., Variable constraints). This design ensures a balanced assessment of code generation, emphasizing both high-level logical coherence and precision in variable management.
3 Metrics
Ensuring that large language models (LLMs) accurately follow instructions is crucial for code generation. To precisely evaluate this capability, we introduce four novel metrics designed to assess how LLMs handle code generation tasks with multiple constraints: Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). These metrics provide a comprehensive evaluation from different perspectives.
For a dataset with $m$ problems, each problem contains a set of $n_{i}$ constraints. We define CSR and SSR as follows:
Completely Satisfaction Rate (CSR)
$\text{CSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\prod_{j=1}^{n_{i}}r_{i,j}\right)$ (1)
where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied. CSR measures the proportion of problems where all constraints are fully met.
Soft Satisfaction Rate (SSR)
$\text{SSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}r_{i,j}}{n_{i}}\right)$ (2)
SSR evaluates the average proportion of constraints satisfied per problem, providing a more flexible assessment.
Rigorous Satisfaction Rate (RSR)
In code generation, some constraints depend on prior instructions, particularly in Combination constraints. To account for dependencies, we define RSR as:
$\text{RSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}\left[r_{i,j}\cdot\prod_{k\in D_{i,j}}r_{i,k}\right]}{n_{i}}\right)$ (3)
where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on. RSR ensures that models satisfy prerequisite constraints before fulfilling dependent ones.
Consistent Continuity Satisfaction Rate (CCSR)
In many code generation tasks, maintaining continuous adherence to instructions is essential. To measure this ability, we define CCSR as:
$\small\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$. CCSR evaluates a model’s consistency in following sequential instructions without errors.
4 Experiment
4.1 Baselines
We evaluate over 30 language models, including both open-source models and commercial API services.
  • Meta Llama 3 Series (Touvron et al., 2023): Includes Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct.
  • Qwen2.5 Series (Yang et al., 2024): Includes Qwen2.5-1.5B-Instruct, Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct. Additionally, it includes dedicated code generation models: Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024), Qwen2.5-Coder-3B-Instruct, Qwen2.5-Coder-7B-Instruct, Qwen2.5-Coder-14B-Instruct, and Qwen2.5-Coder-32B-Instruct.
  • Mistral Series (Jiang et al., 2023): Includes Mistral-3B, Mistral-8B, and the specialized code model Codestral-2501.
  • Microsoft Phi Series (Abdin et al., 2024): Includes Phi-3.5-Mini-128K-Instruct (3.8B) and the latest Phi-4.
  • Google Gemma Series (Team, 2024b): Includes Gemma-2-9B-It and Gemma-2-27B-It.
  • DeepSeek Series: Includes DeepSeek-Coder (Guo et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024).
  • OpenAI Series (Achiam et al., 2023): Includes GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and the latest GPT-4O-2024-11-20.
  • Google Gemini Series (Team, 2024a): Includes Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro.
  • Anthropic Series: Includes Claude-3.5-Sonnet-20241022.
4.2 Experimental Setup
The temperature coefficient is set to 0 to ensure output determinism, with a maximum generation length of 4096 tokens. All other settings follow the official default parameters for each model. Commercial API models are accessed through the latest available interface as of December 2024. Specifically, this study evaluates the Qwen2.5-Coder series, Meta’s latest Llama 3.3 architecture, and the DeepSeek-V3 large-scale dialogue model as upper-bound benchmarks. All experiments are conducted using the official API and 8 H800(80G).
4.3 Automatic Evaluation
To ensure robust evaluation, we employed both large language models (LLMs) and human experts to assess whether each model correctly followed the atomic constraint instructions. During data preprocessing, all constraints were decomposed into atomic elements, allowing the adjudication system to perform objective binary evaluations (Yes/No) instead of relying on subjective judgment.
Follow by FairEval (Wang et al., 2023a), GPT-4-1106-Preview model served as the primary automated evaluation tool (see Appendix 8.1 for the detailed prompt). Additionally, three domain experts manually annotated a stratified sample of 150 instances. Statistical analysis revealed strong agreement between LLM-based and human evaluations, with Pearson correlation coefficients of 0.87 (LLM-human) and 0.85 (inter-human), demonstrating high consistency between automated and manual assessments.
These results suggest that the atomic constraint decomposition methodology effectively reduces subjectivity, making LLM-based evaluation a reliable and scalable approach for constraint verification tasks.
4.4 Main Experiments
Table 3: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
Models CSR SSR RSR CCSR
Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
Llama-3.2-1B-Instruct 0.034 0.046 0.022 0.218 0.277 0.159 0.182 0.231 0.132 0.152 0.197 0.107
Llama-3.1-8B-Instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.340 0.370 0.444 0.295
Qwen2.5-Coder-7B-Instruct 0.142 0.198 0.087 0.514 0.590 0.438 0.453 0.533 0.373 0.390 0.463 0.318
Qwen2.5-7B-Instruct 0.153 0.201 0.104 0.535 0.599 0.471 0.475 0.546 0.405 0.416 0.479 0.353
Ministral-8B 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.490 0.371
Gemma-2-9B-It 0.171 0.210 0.131 0.573 0.642 0.504 0.513 0.587 0.440 0.445 0.508 0.383
Qwen2.5-Coder-32B-Instruct 0.365 0.422 0.307 0.736 0.767 0.704 0.679 0.723 0.635 0.634 0.669 0.599
Gemma-2-27B-It 0.245 0.300 0.190 0.658 0.709 0.607 0.596 0.652 0.540 0.533 0.588 0.478
Qwen2.5-32B-Instruct 0.294 0.337 0.251 0.680 0.722 0.638 0.621 0.674 0.568 0.560 0.604 0.515
Qwen2.5-72B-Instruct 0.281 0.319 0.244 0.685 0.734 0.634 0.621 0.677 0.564 0.569 0.619 0.518
Llama-3.3-70B-Instruct 0.307 0.359 0.255 0.698 0.749 0.647 0.632 0.691 0.574 0.589 0.643 0.536
Gemini-Exp-1206 0.357 0.410 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
GPT-4O-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.650 0.698 0.602
Claude-3-5-Sonnet-20241022 0.444 0.525 0.362 0.727 0.784 0.669 0.692 0.757 0.626 0.652 0.715 0.587
Deepseek-V3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.680
Table 3 provides a comprehensive evaluation of CodeIF. The results are measured using four key metrics: Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR). More detailed results are in Appendix 8.2.
Result Overview
DeepSeek-V3 and Claude-3-5-Sonnet-20241022 consistently achieve top scores across all difficulty levels and evaluation metrics, excelling in complex tasks. However, even the highest CSR on the Hard dataset is only 0.362, highlighting room for improvement in handling strict constraints.
Model Scale and Performance Trends
Larger models generally perform better across all metrics. In the Qwen2.5 series (Qwen2.5-1.5B-Instruct to Qwen2.5-72B-Instruct), performance improves with model size. However, this trend is not universal. In the Llama3 series, larger models do not always achieve the best results, suggesting that architecture design, training data quality, and optimization strategies are equally critical factors.
Open-Source vs. Closed-Source Models
Closed-source models like GPT-4O and Claude-3-5-Sonnet-20241022 outperform their open-source counterparts, maintaining higher accuracy under complex constraints. While some large open-source models (e.g., Qwen2.5-72B-Instruct) show competitive results, they still lag behind. This performance gap likely stems from advantages in data quality, optimization techniques, and reinforcement learning from human feedback (RLHF), which proprietary models leverage for enhanced instruction-following precision.
Impact of Task Difficulty
Table 3 shows that as task complexity increases (from Easy to Hard), the performance of nearly all models declines. For example, GPT-4O-2024-11-20 achieves a CSR of 0.441 on Easy tasks, but this drops to 0.325 on Hard tasks. This trend confirms that higher constraint levels impose greater challenges, reducing instruction-following accuracy. These findings align with previous research, which suggests that models’ ability to adhere to instructions deteriorates as task difficulty increases.
5 In-Depth Analysis
Figure 7: Performance of different LLMs on the CodeIF evaluation across instruction categories, measured by SSR.
5.1 Performance Analysis Across Instruction Types
Figure 7 illustrates the performance of LLMs across different instruction categories, revealing notable disparities among models. DeepSeek-V3 consistently achieves the highest scores across most evaluation dimensions, with a particularly strong performance in combination tasks (0.831), highlighting its effectiveness in handling complex instructions. However, its strengths are not uniform; it excels in global structure control but shows relative weaknesses in variable handling, likely reflecting its specific optimization strategies.
Conversely, Meta’s Llama series demonstrates a significant correlation between model size and performance. While smaller configurations (Llama-3.2-1B-Instruct) score lower, larger variants (Llama-3.3-70B-Instruct) perform significantly better, showcasing substantial gains with increased parameter counts. However, model size alone is not the sole determinant of performance. Comparisons between similarly sized models, such as Google’s Gemma and Meta’s Llama, indicate distinct strengths in specific metrics, suggesting that architectural design and training methodologies also play a crucial role.
5.2 Cross-Language Performance Analysis of LLMs
Figure 8: SSR scores of LLMs across different programming languages in the CodeIF evaluation.
Figure 8 compares the performance of leading LLMs across four programming languages: C++, Java, Python, and Go, highlighting key trends at both the model and language levels.
At the model level, DeepSeek-V3 consistently outperforms others, achieving the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), with an RSR of 0.787 in Java, likely benefiting from pretraining optimizations for code continuity. Claude-3-5-Sonnet excels in Java, attaining the highest CSR (0.504) and RSR (0.749), indicating specialized mechanisms for handling complex constraints. However, its SSR in Python (0.703) is relatively lower, suggesting challenges in cross-language generalization. GPT-4O demonstrates balanced performance across languages, ranking second in CSR (0.355) and RSR (0.682) for Python, with minimal variation across different programming paradigms (coefficient of variation CV = 0.18).
At the language level, C++ presents the most significant challenges, likely due to its complex template metaprogramming, which increases constraint difficulty. Java exhibits the highest inter-model variance, with Claude-3-5-Sonnet standing out, possibly due to optimizations tailored to its strong type system. Go achieves the highest overall SSR, benefiting from its simple syntax, but RSR fluctuates significantly, reflecting differences in how models handle concurrency constraints.
These findings highlight variations in LLMs’ ability to generalize across languages and suggest potential areas for optimization, such as improving handling of dependency constraints and enhancing consistency across different programming paradigms.
5.3 Analysis of Instruction Adherence Deviations
Analysis of model-generated responses reveals frequent deviations from specific instructions, particularly in naming conventions and formatting constraints. Models often overlook global formatting rules, such as line limits and character restrictions, when handling general structure instructions.
Moreover, adherence to naming conventions remains inconsistent. For instance, even when explicitly instructed to use PascalCase for variable names, models frequently generate results in lowercase or underscore-separated formats (e.g., transforming current_power into CurrentPower incorrectly).
A particularly notable issue is the tendency to disregard prohibitive instructions. For example, when explicitly instructed to avoid if statements in decision-making—favoring ternary operators or data structures (e.g., dictionaries)—models still generate outputs containing if statements, indicating gaps in constraint enforcement.
5.4 Enhance Model Instruction Compliance
Insights from Appendix Table 5 highlight key strategies for improving instruction adherence in code generation. Supervised Fine-Tuning (SFT) has proven effective, particularly in the Llama series, while model scale remains a crucial factor—larger models like Qwen2.5-72B-Instruct consistently outperform smaller counterparts such as Qwen2.5-3B-Instruct in instruction-following accuracy.
To enhance adherence, several approaches can be employed. First, classifying instructions by priority ensures strict enforcement of hard constraints (e.g., syntax restrictions) while allowing flexibility for soft guidelines (e.g., coding styles). Patterned code generation can provide structured alternatives, such as replacing conditional statements with lookup tables or state machines. Integrating a naming convention engine allows automated conversion of variable and class names (e.g., transforming snake_case to PascalCase). Additionally, leveraging Abstract Syntax Tree (AST) analysis helps detect and transform prohibited structures, such as replacing for loops with while loops or recursive calls. Finally, conflict resolution mechanisms can be introduced to handle contradictory user instructions, such as providing alternative solutions when certain language features are unavailable (e.g., requesting switch-case in Python).
6 Related Work
Code generation and instruction-following are pivotal capacities under examination in AI research (Feng et al., 2020; Sun et al., 2024; Luo et al., 2024; Wang et al., 2023b; Kim et al., 2018; Li et al., 2023; Lu et al., 2021; Li et al., 2022; Wei et al., 2023; Nijkamp et al., 2023b; Zhuo et al., 2024; Jain et al., 2024; Nijkamp et al., 2023a; Zhang et al., 2023; Allal et al., 2023; Lozhkov et al., 2024a; Roziere et al., 2023; Lozhkov et al., 2024b; Wang et al., 2021; Yan et al., 2023). Several benchmarks have been devised to appraise these capabilities in large-scale models. For code generation, benchmarks like McEval (Chai et al., 2024), FullStackBench (Liu et al., 2024), Repocoder (Zhang et al., 2023), Repobench (Liu et al., 2023), and LiveCodeBench (Jain et al., 2024) have been notable. Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench (Qin et al., 2024), CFBench (Zhang et al., 2024), Instruct-following (Zhou et al., 2023), and FollowBench (Jiang et al., 2024), each tailored to assess different aspects of following instructions given to models.
7 Conclusion
This study introduces CodeIF, a benchmark for evaluating the instruction-following capabilities of LLMs in code generation. Covering Java, Python, Go, and C++, CodeIF constructs a diverse test set with constraints ranging from global to specific variables. It introduces novel evaluation metrics—Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR)—to assess multi-constraint handling across multiple dimensions.

### v3

@@
 [^1] CodeIF data and code are publicly available: https://github.com/lin-rany/codeIF.
-Machine Learning, ICML
-1Beihang University, 2Tsinghua University, 3National University of Singapore
+CodeIF: Benchmarking the Instruction-Following Capabilities of Large Language Models for Code Generation
+Kaiwen Yan1 *, Hongcheng Guo1 * †, Xuanqing Shi2 Shaosheng Cao3, Donglin Di2, Zhoujun Li1 1Beihang University, 2Tsinghua University, 3Xiaohongshu lin_rany@foxmail.com, hongchengguo@buaa.edu.cn, sxq23@mails.tsinghua.edu.cn caoshaosheng@xiaohongshu.com, donglin.ddl@gmail.com, lizj@buaa.edu.cn
 footnotetext: *Equal contributionfootnotetext: Corresponding Author
@@
 With the rapid advancement of large language models (LLMs), automated code generation is undergoing a profound transformation.
-While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge (Liu et al., 2024; Zhang et al., 2023).
-To drive progress in this field, a comprehensive and systematic evaluation framework is essential (Jiang et al., 2024; Zhou et al., 2023).
+While LLMs have demonstrated promising capabilities in programming tasks, their ability to comprehend and execute complex instructions remains a challenge Liu et al.
+(2024); Zhang et al.
+(2023).
+To drive progress in this field, a comprehensive and systematic evaluation framework is essential Jiang et al.
+(2024); Zhou et al.
+(2023).
 This study introduces CodeIF, a benchmark designed to assess LLMs’ instruction-following capabilities in code generation.
-Built upon insights from existing evaluation sets like McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++.
+Built upon insights from existing evaluation sets like McEval Chai et al.
+(2024) and FullStackBench Liu et al.
+(2024), CodeIF is tailored for multi-language environments, covering Java, Python, Go, and C++.
 It categorizes tasks by difficulty and systematically evaluates models across 50 fine-grained sub-instructions, providing a nuanced understanding of their strengths and weaknesses.
@@
   1. 1.
-    Innovative Benchmark: We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation.
+    Innovative Benchmark.
+We introduce CodeIF, the first systematic benchmark for evaluating LLMs’ instruction-following capabilities in code generation.
 CodeIF categorizes tasks into 8 main types and 50 fine-grained sub-instructions, ensuring a comprehensive assessment of model performance.
   2. 2.
-    Automated High-Quality Instruction Generation: Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists.
+    Automated High-Quality Instruction Generation.
+Leveraging advanced LLMs like GPT-4, we develop a method to automatically generate constraint-based instruction lists.
 This approach enhances evaluation depth by incorporating instructional dependencies while minimizing human intervention.
   3. 3.
-    Novel Evaluation Metrics: We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks.
+    Novel Evaluation Metrics.
+We propose a new framework with four key metrics (CSR, SSR, RSR, and CCSR) tailored for code generation tasks.
 These metrics assess models’ ability to handle multi-constraint problems across different dimensions, offering deeper insights and new benchmarks for future research.
   4. 4.
-    Extensive Evaluation and Analysis: We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels.
+    Extensive Evaluation and Analysis.
+We systematically evaluate 35 state-of-the-art LLMs, including both open-source and commercial models, across multiple programming languages and difficulty levels.
 Our experiments uncover current strengths and limitations, providing clear directions for future advancements.
@@
 The first step involves the construction of constraint instructions, followed by the assembly of the dataset, and finally the construction of dependencies between instructions.
-As shown in Figure 1, the construction of CodeIF includes the following steps: we first collect and refine constraint instructions from real code generation tasks, then we combine code generation tasks with LLMs and human review to create a high-quality evaluation dataset.
+As shown in Figure 1, CodeIF is constructed by collecting and refining constraint instructions from real code generation tasks, then combining these tasks with LLM outputs and human review to create a high-quality evaluation dataset.
 2.1 Building
-The construction of the CodeIF dataset consists of two main phases: collection of constraint instructions and data processing.
-In Section  2.2, we elaborate on how we collected the constraint instructions for code generation.
-Following this, in Section  2.3, we describe how we combined these constraint instructions to produce the final CodeIF evaluation dataset.
+The construction of the CodeIF dataset involves two phases: collecting constraint instructions (Section 2.2) and processing data to create the final CodeIF evaluation dataset (Section 2.3).
 2.2 Constraint Instructions Collection
-The first phase of our work focuses on code generation by collecting and formulating constraint instructions to construct the CodeIF evaluation dataset.
-This process consists of two stages: (1) collecting and verifying constraint instructions and (2) applying them to dataset generation.
-We begin with an in-depth analysis of code generation tasks based on existing benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024), establishing an instruction system covering eight categories.
-Each category targets a specific aspect of code generation, enabling a fine-grained assessment of LLMs’ instruction-following abilities.
-Within each category, we further refine constraints into atomic instructions with clear, explicit directives.
-These atomic constraints allow binary evaluation (“yes” or “no”), ensuring an objective and precise assessment of model compliance while minimizing subjectivity.
-Our classification of atomic constraint instructions is guided by an analysis of existing code generation benchmarks and widely used code repositories.
-We observe that requirement distribution varies significantly across different levels—line-level, function-level, and full-solution generation.
-To ensure comprehensive coverage, we design atomic instruction categories that assess both high-level architectural constraints and fine-grained variable control.
-To systematically evaluate LLMs’ instruction-following capabilities in code generation, we define eight constraint categories.
-The Global category ensures that generated code aligns with overarching specifications and holistic constraints.
-Structural Control focuses on adherence to control structures such as loops and conditionals, as well as data structures like sets and maps.
-Variable constraints assess details at the variable level, including naming conventions and initialization methods.
-At different abstraction levels, Interface, Function, and Class constraints evaluate the model’s ability to construct well-defined program components.
-The File category tests handling of cross-file dependencies and external library calls.
-Finally, the Combination category presents a comprehensive challenge by integrating constraints from multiple dimensions, assessing how well models follow complex, multi-faceted instructions in realistic scenarios.
+The first phase of our work centers on code generation, constructing the CodeIF evaluation dataset through two steps: (1) collecting and verifying constraint instructions, and (2) using them for dataset generation.
+We analyze benchmarks like McEval Chai et al.
+(2024) and FullStackBench Liu et al.
+(2024) to develop an instruction system spanning eight categories, each targeting specific aspects of code generation for a fine-grained assessment of LLMs’ instruction-following abilities.
+Constraints are decomposed into atomic instructions with explicit directives, enabling objective binary evaluation (yes/no) and minimizing subjective interpretation.
+The eight categories cover both architectural-level specifications and variable-level implementation controls, ensuring comprehensive constraint coverage.
+Specifically, the Global category evaluates adherence to overarching specifications, while Structural Control focuses on control structures (e.g., loops, conditionals) and data structures.
+Variable constraints assess naming and initialization.
+Higher abstraction levels include Interface, Function, and Class constraints for program components, while the File category tests cross-file dependencies and external library handling.
+The Combination category integrates constraints across dimensions, challenging models with complex scenarios.
 Figure 2: CodeIF Constraints Instruction Distribution
-As shown in Figure 2, CodeIF’s distribution across different programming languages and categories is presented.
-Through a detailed analysis of existing code generation benchmarks, we designed an evaluation system comprising 8 main categories and 50 fine-grained constraint instructions.
-This system not only enables a systematic assessment of LLMs’ performance in code generation but also establishes clear benchmarks for future research.
-These instructions cover constraints ranging from high-level structural guidelines to fine-grained variable operations, ensuring a comprehensive evaluation.
-For instance, the Global category assesses a model’s ability to maintain overall structural integrity and logical coherence, while the Variable category evaluates its precision in variable naming and initialization.
-By refining constraints into well-defined categories, we can precisely identify the strengths and weaknesses of different LLMs, guiding optimization efforts and advancing the field toward more effective automated code generation.
-Additionally, this structured evaluation approach provides a valuable reference for future research, fostering continuous innovation.
-The complete list of atomic constraint instructions is provided in Appendix Table 4.
+Figure 2 presents CodeIF’s distribution across programming languages and categories.
+The evaluation system features 8 categories and 50 fine-grained constraint instructions, systematically assessing LLMs’ code generation performance.
+By organizing constraints clearly, the system identifies strengths and weaknesses, guides optimization, and advances automated code generation.
+The full list of constraints is in Appendix 4.
 2.3 Data Construction
 Multi-Language and Difficulty-Differentiated Benchmark Design
-To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval (Chai et al., 2024) and FullStackBench (Liu et al., 2024).
+To ensure diversity and comprehensiveness in evaluation, we carefully selected code generation tasks across four mainstream programming languages—Java, Python, Go, and C++—from leading benchmarks such as McEval Chai et al.
+(2024) and FullStackBench Liu et al.
+(2024).
 These languages, spanning both dynamic and static paradigms, create a rich linguistic environment that enhances multi-language assessment.
@@
 Automated Generation of Constraint Instructions
-To construct high-quality constraint instructions, we leveraged large language models (LLMs) such as GPT-4 to automatically generate task-specific instruction lists.
-Initially, we curated approximately 20 detailed examples, each featuring a specific code generation task.
-Human experts then formulated atomic constraint instructions to ensure precise adherence during code generation.
-These examples served as prompts, allowing LLMs to generate corresponding instruction lists while refining task descriptions by identifying and eliminating redundant constraints.
-This approach streamlined instruction-following requirements, ensuring clarity and promoting high-quality outputs.
+We used large language models (LLMs) like GPT-4 to create task-specific instruction lists for code generation tasks.
+We prepared 20 detailed examples and formulated concise atomic instructions for accuracy.
+These examples guided LLMs in refining tasks and streamlining instructions to enhance clarity and output quality.
 Constructing Instruction Dependencies
-Since some atomic constraints exhibit interdependencies, we further utilized LLMs to construct explicit dependency relationships within instruction sets.
-These dependencies were integrated into our evaluation framework, enhancing assessment depth and improving the accuracy and practicality of constraint verification.
-Specifically, by analyzing the logical connections and sequential dependencies among atomic instructions, we can precisely define the execution steps required for each task.
-For instance, in function generation tasks, the process typically involves defining the function name first, specifying parameter types, and finally implementing the function body.
-These dependencies ensure coherence and consistency throughout the code generation process.
-Integrating these dependencies into the evaluation system enables a more accurate assessment of a model’s ability to follow complex instructions while providing a foundation for future optimizations.
-Additionally, this approach helps identify potential challenges models face when handling specific instruction sequences, offering insights for further improvement.
-As shown in Figure 3, this example from CodeIF illustrates a specific generation task, its associated instruction list, and the corresponding instruction dependencies.
+We utilized LLMs to map dependencies between atomic constraints, improving our evaluation framework’s precision and verification accuracy.
+By understanding the dependencies among instructions, we outlined clear steps for tasks like function creation, which involve naming the function, defining parameter types, and coding the body.
+Incorporating these dependencies enhances our evaluation system, more accurately assessing the model’s capability with complex instructions and identifying areas for improvement.
+Figure 3 illustrates a CodeIF task with its instruction sequence and dependencies.
 Figure 3: Specific cases of the CodeIF dataset, ’Task’ denotes the specific generation task, ’Type’ refers to the type of constraint, and ’Dependence’ indicates the prerequisite constraints for this constraint.
 2.4 Data Analysis
-Table 1: CodeIF dataset statistics, showing the statistical information of different difficulty classifications. Avg.Instr represents the average length of the atomic constraint instruction list.
 Set Num Avg.Instr Go Python Java C++
@@
 Full 1200 12.90 230 348 353 269
+Table 1: CodeIF dataset statistics, showing the statistical information of different difficulty classifications. Avg.Instr represents the average length of the atomic constraint instruction list.
 CodeIF Static Analysis
 Figure 4: The distribution of atomic instruction list lengths in datasets of different difficulties.
-As shown in Table 1, the dataset is categorized into three difficulty levels: Easy, Hard, and an aggregated set, Full.
-The Easy and Hard sets each contain 600 code generation tasks, while the Full dataset combines both, totaling 1,200 tasks.
-These tasks span four programming languages: Go, Python, Java, and C++.
-Among them, Java has the highest number of tasks (353), followed by Python (348).
-C++ and Go contain 269 and 230 tasks, respectively.
-The dataset also records the average number of instructions per task for each difficulty level.
-The Easy set averages 11.99 instructions per task, while the Hard set increases to 13.8, reflecting greater complexity and stricter constraints.
-The Full dataset, combining both, has an average instruction length of 12.9.
-Figure 4 illustrates the distribution of task lengths across the dataset.
-Table 2: CodeIF dataset statistics information, showing the distribution of atomic restriction instruction categories under different difficulty classifications.
+Table 1 categorizes the dataset into three difficulty levels: Easy, Hard, and Full.
+Both Easy and Hard sets contain 600 tasks, while the Full dataset combines them, totaling 1,200 tasks across Go, Python, Java, and C++.
+Java has the most tasks (353), followed by Python (348), C++ (269), and Go (230).
+The Easy set averages 11.99 instructions per task, the Hard set 13.8, and the Full dataset 12.9, reflecting increasing complexity.
+Figure 4 shows task length distribution.
 Set Global Structural Control Variable Interface Function Class File Combination
@@
 Full 3528 2201 2815 932 1228 1167 1525 2095
-Figure 5: The distribution of constraint instruction list lengths in datasets of different difficulties.
-Analysis of Constraint Instruction Category Distribution
-Table 2 presents the distribution of atomic instructions across different difficulty levels in the evaluation set.
-As shown in Figure 5, we compare the average number of instructions per category between the “Easy” and “Hard” sets.
-These categories include Global, Structural Control, Variable, Interface, Function, Class, File, and Combination.
-Across all categories, the “Hard” dataset contains a higher average number of instructions than the “Easy” set.
-Notably, the Global category has the highest instruction count in both difficulty levels, averaging 2.5 instructions in the “Easy” set and exceeding 3 in the “Hard” set.
-This trend indicates that as task complexity increases, models face greater challenges in handling intricate constraints and detailed requirements.
-By analyzing instruction distribution across categories, we gain deeper insights into model performance under varying levels of difficulty, providing a foundation for further optimization.
-Figure 6: Distribution of atomic instruction list lengths across difficulty levels.
-Figure 6 illustrates the proportion of each atomic instruction category.
-The Global category holds the largest share at 22.77%, followed by the Variable category at 18.17%.
-This distribution highlights CodeIF’s focus on evaluating LLMs’ capabilities at both the overall structural level (e.g., Global constraints) and fine-grained variable handling (e.g., Variable constraints).
-This design ensures a balanced assessment of code generation, emphasizing both high-level logical coherence and precision in variable management.
+Table 2: CodeIF dataset statistics information, showing the distribution of atomic restriction instruction categories under different difficulty classifications.
+Constraint Instruction Analysis
+Table 2compares instruction distribution across difficulty levels.
+The Hard set consistently has more instructions per category than the Easy set, with the Global category averaging 2.5 instructions in Easy and over 3 in Hard.
+This indicates greater challenges for models as task complexity rises.
+More analysis is in Appendix D.
 3 Metrics
@@
 For a dataset with $m$ problems, each problem contains a set of $n_{i}$ constraints. We define CSR and SSR as follows:
-Completely Satisfaction Rate (CSR)
+Completely Satisfaction Rate (CSR):
 $\text{CSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\prod_{j=1}^{n_{i}}r_{i,j}\right)$ (1)
-where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied. CSR measures the proportion of problems where all constraints are fully met.
-Soft Satisfaction Rate (SSR)
+where $r_{i,j}\in[0,1]$ indicates whether the $j$-th constraint in the $i$-th problem is satisfied.
+Soft Satisfaction Rate (SSR):
 $\text{SSR}=\frac{1}{m}\sum_{i=1}^{m}\left(\frac{\sum_{j=1}^{n_{i}}r_{i,j}}{n_{i}}\right)$ (2)
@@
 where $D_{i,j}$ represents the set of constraints that the $j$-th constraint in the $i$-th problem depends on.
-RSR ensures that models satisfy prerequisite constraints before fulfilling dependent ones.
 Consistent Continuity Satisfaction Rate (CCSR)
 In many code generation tasks, maintaining continuous adherence to instructions is essential. To measure this ability, we define CCSR as:
-$\small\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
-where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$. CCSR evaluates a model’s consistency in following sequential instructions without errors.
+$\text{CCSR}=\frac{1}{m}\sum_{i=1}^{m}\frac{L_{i}}{n_{i}},\\ L_{i}=\max\Bigl{\{}l\,\Big{|}\,\exists t\mathbin{\in}[1,n_{i}{-}l{+}1],\ \prod_{\mathclap{j=t}}^{\mathclap{t+l-1}}r_{i,j}=1\Bigr{\}}$ (4)
+where $L_{i}$ represents the longest consecutive sequence of satisfied constraints in problem $i$.
 4 Experiment
-4.1 Baselines
-We evaluate over 30 language models, including both open-source models and commercial API services.
-  • Meta Llama 3 Series (Touvron et al., 2023): Includes Llama-3.2-1B-Instruct, Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, and Llama-3.3-70B-Instruct.
-  • Qwen2.5 Series (Yang et al., 2024): Includes Qwen2.5-1.5B-Instruct, Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, and Qwen2.5-72B-Instruct.
-Additionally, it includes dedicated code generation models: Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024), Qwen2.5-Coder-3B-Instruct, Qwen2.5-Coder-7B-Instruct, Qwen2.5-Coder-14B-Instruct, and Qwen2.5-Coder-32B-Instruct.
-  • Mistral Series (Jiang et al., 2023): Includes Mistral-3B, Mistral-8B, and the specialized code model Codestral-2501.
-  • Microsoft Phi Series (Abdin et al., 2024): Includes Phi-3.5-Mini-128K-Instruct (3.8B) and the latest Phi-4.
-  • Google Gemma Series (Team, 2024b): Includes Gemma-2-9B-It and Gemma-2-27B-It.
-  • DeepSeek Series: Includes DeepSeek-Coder (Guo et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024).
-  • OpenAI Series (Achiam et al., 2023): Includes GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and the latest GPT-4O-2024-11-20.
-  • Google Gemini Series (Team, 2024a): Includes Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro.
-  • Anthropic Series: Includes Claude-3.5-Sonnet-20241022.
 4.1 Experimental Setup
@@
 Commercial API models are accessed through the latest available interface as of December 2024.
-Specifically, this study evaluates the Qwen2.5-Coder series, Meta’s latest Llama 3.3 architecture, and the DeepSeek-V3 large-scale dialogue model as upper-bound benchmarks.
 All experiments are conducted using the official API and 8 H800(80G).
 4.2 Automatic Evaluation
-To ensure robust evaluation, we employed both large language models (LLMs) and human experts to assess whether each model correctly followed the atomic constraint instructions.
-During data preprocessing, all constraints were decomposed into atomic elements, allowing the adjudication system to perform objective binary evaluations (Yes/No) instead of relying on subjective judgment.
-Follow by FairEval (Wang et al., 2023a), GPT-4-1106-Preview model served as the primary automated evaluation tool (see Appendix 8.1 for the detailed prompt).
-Additionally, three domain experts manually annotated a stratified sample of 150 instances.
-Statistical analysis revealed strong agreement between LLM-based and human evaluations, with Pearson correlation coefficients of 0.87 (LLM-human) and 0.85 (inter-human), demonstrating high consistency between automated and manual assessments.
-These results suggest that the atomic constraint decomposition methodology effectively reduces subjectivity, making LLM-based evaluation a reliable and scalable approach for constraint verification tasks.
+To ensure robust evaluation, we used LLMs and human experts to verify model adherence to atomic constraints.
+Constraints were decomposed into atomic elements, enabling objective binary evaluations (Yes/No) over subjective judgments.
+Following FairEval Wang et al.
+(2023a), GPT-4-1106-Preview was the primary evaluation tool (prompt details in Appendix A).
+Three domain experts manually annotated 150 stratified samples.
+Statistical analysis showed strong agreement, with Pearson correlations of 0.87 (LLM-human) and 0.85 (inter-human), confirming high consistency.
+Baselines are in Appendix C.
 4.3 Main Experiments
-Table 3: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
 Models CSR SSR RSR CCSR
@@
 Deepseek-V3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.680
-Table 3 provides a comprehensive evaluation of CodeIF.
-The results are measured using four key metrics: Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR).
-More detailed results are in Appendix 8.2.
-Result Overview
-DeepSeek-V3 and Claude-3-5-Sonnet-20241022 consistently achieve top scores across all difficulty levels and evaluation metrics, excelling in complex tasks.
-However, even the highest CSR on the Hard dataset is only 0.362, highlighting room for improvement in handling strict constraints.
-Model Scale and Performance Trends
-Larger models generally perform better across all metrics.
-In the Qwen2.5 series (Qwen2.5-1.5B-Instruct to Qwen2.5-72B-Instruct), performance improves with model size.
-However, this trend is not universal.
-In the Llama3 series, larger models do not always achieve the best results, suggesting that architecture design, training data quality, and optimization strategies are equally critical factors.
-Open-Source vs. Closed-Source Models
-Closed-source models like GPT-4O and Claude-3-5-Sonnet-20241022 outperform their open-source counterparts, maintaining higher accuracy under complex constraints.
-While some large open-source models (e.g., Qwen2.5-72B-Instruct) show competitive results, they still lag behind.
-This performance gap likely stems from advantages in data quality, optimization techniques, and reinforcement learning from human feedback (RLHF), which proprietary models leverage for enhanced instruction-following precision.
-Impact of Task Difficulty
-Table 3 shows that as task complexity increases (from Easy to Hard), the performance of nearly all models declines.
-For example, GPT-4O-2024-11-20 achieves a CSR of 0.441 on Easy tasks, but this drops to 0.325 on Hard tasks.
-This trend confirms that higher constraint levels impose greater challenges, reducing instruction-following accuracy.
-These findings align with previous research, which suggests that models’ ability to adhere to instructions deteriorates as task difficulty increases.
+Table 3: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
+Table 3 evaluates CodeIF using four metrics: CSR, SSR, RSR, and CCSR. Detailed results are in Appendix B.
+Overview.
+DeepSeek-V3 and Claude-3-5-Sonnet-20241022 lead across metrics, excelling in complex tasks. However, the highest CSR on Hard tasks is just 0.362, showing challenges in meeting strict constraints.
+Model Scale Trends.
+Larger models generally perform better, as seen in Qwen2.5 series. However, the Llama3 series shows inconsistent results, highlighting the importance of architecture, data quality, and optimization.
+Open vs. Closed Models.
+Closed-source models like GPT-4O and Claude-3-5 outperform open-source models, especially under complex constraints.
+While large open-source models (e.g., Qwen2.5-72B-Instruct) are competitive, they lag due to differences in data quality and RLHF techniques.
+Task Difficulty Impact.
+Performance drops with increasing task complexity. For instance, GPT-4O’s CSR falls from 0.441 on Easy tasks to 0.325 on Hard tasks, highlighting the challenge of strict constraints.
 5 In-Depth Analysis
@@
 5.1 Performance Analysis Across Instruction Types
-Figure 7 illustrates the performance of LLMs across different instruction categories, revealing notable disparities among models.
-DeepSeek-V3 consistently achieves the highest scores across most evaluation dimensions, with a particularly strong performance in combination tasks (0.831), highlighting its effectiveness in handling complex instructions.
-However, its strengths are not uniform; it excels in global structure control but shows relative weaknesses in variable handling, likely reflecting its specific optimization strategies.
-Conversely, Meta’s Llama series demonstrates a significant correlation between model size and performance.
-While smaller configurations (Llama-3.2-1B-Instruct) score lower, larger variants (Llama-3.3-70B-Instruct) perform significantly better, showcasing substantial gains with increased parameter counts.
-However, model size alone is not the sole determinant of performance.
-Comparisons between similarly sized models, such as Google’s Gemma and Meta’s Llama, indicate distinct strengths in specific metrics, suggesting that architectural design and training methodologies also play a crucial role.
+Figure 5 compares LLM performance across instruction categories, revealing notable variations.
+DeepSeek-V3 leads overall, excelling in combination tasks (0.831) and global structure control, though weaker in variable handling, reflecting its optimization focus.
+Meta’s Llama series shows a clear correlation between model size and performance, with larger variants (Llama-3.3-70B-Instruct) outperforming smaller ones (Llama-3.2-1B-Instruct).
+However, size alone is not decisive; comparisons with similarly sized models like Google’s Gemma highlight the role of architecture and training methods in shaping performance.
 5.2 Cross-Language Performance Analysis of LLMs
 Figure 6: SSR scores of LLMs across different programming languages in the CodeIF evaluation.
-Figure 8 compares the performance of leading LLMs across four programming languages: C++, Java, Python, and Go, highlighting key trends at both the model and language levels.
-At the model level, DeepSeek-V3 consistently outperforms others, achieving the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), with an RSR of 0.787 in Java, likely benefiting from pretraining optimizations for code continuity.
-Claude-3-5-Sonnet excels in Java, attaining the highest CSR (0.504) and RSR (0.749), indicating specialized mechanisms for handling complex constraints.
-However, its SSR in Python (0.703) is relatively lower, suggesting challenges in cross-language generalization.
-GPT-4O demonstrates balanced performance across languages, ranking second in CSR (0.355) and RSR (0.682) for Python, with minimal variation across different programming paradigms (coefficient of variation CV = 0.18).
-At the language level, C++ presents the most significant challenges, likely due to its complex template metaprogramming, which increases constraint difficulty.
-Java exhibits the highest inter-model variance, with Claude-3-5-Sonnet standing out, possibly due to optimizations tailored to its strong type system.
-Go achieves the highest overall SSR, benefiting from its simple syntax, but RSR fluctuates significantly, reflecting differences in how models handle concurrency constraints.
-These findings highlight variations in LLMs’ ability to generalize across languages and suggest potential areas for optimization, such as improving handling of dependency constraints and enhancing consistency across different programming paradigms.
+Figure 6 compares the performance of leading LLMs across C++, Java, Python, and Go, highlighting trends at both model and language levels.
+At the model level, DeepSeek-V3 leads with the highest CCSR in C++ (0.725), Java (0.753), and Go (0.722), and an RSR of 0.787 in Java.
+Claude-3-5-Sonnet excels in Java with the highest CSR (0.504) and RSR (0.749), but shows lower SSR in Python (0.703).
+GPT-4O demonstrates balanced performance, ranking second in Python’s CSR (0.355) and RSR (0.682), with minimal variance (CV = 0.18).
+At the language level, C++ is the most challenging due to complex template metaprogramming.
+Java shows high inter-model variance, with Claude-3-5-Sonnet performing best.
+Go achieves the highest SSR but fluctuates in RSR.
+These results highlight cross-language generalization differences and suggest optimization areas like dependency handling and paradigm consistency.
 5.3 Analysis of Instruction Adherence Deviations
-Analysis of model-generated responses reveals frequent deviations from specific instructions, particularly in naming conventions and formatting constraints.
-Models often overlook global formatting rules, such as line limits and character restrictions, when handling general structure instructions.
-Moreover, adherence to naming conventions remains inconsistent.
-For instance, even when explicitly instructed to use PascalCase for variable names, models frequently generate results in lowercase or underscore-separated formats (e.g., transforming current_power into CurrentPower incorrectly).
-A particularly notable issue is the tendency to disregard prohibitive instructions.
-For example, when explicitly instructed to avoid if statements in decision-making—favoring ternary operators or data structures (e.g., dictionaries)—models still generate outputs containing if statements, indicating gaps in constraint enforcement.
-5.4 Enhance Model Instruction Compliance
-Insights from Appendix Table 5 highlight key strategies for improving instruction adherence in code generation.
-Supervised Fine-Tuning (SFT) has proven effective, particularly in the Llama series, while model scale remains a crucial factor—larger models like Qwen2.5-72B-Instruct consistently outperform smaller counterparts such as Qwen2.5-3B-Instruct in instruction-following accuracy.
-To enhance adherence, several approaches can be employed.
-First, classifying instructions by priority ensures strict enforcement of hard constraints (e.g., syntax restrictions) while allowing flexibility for soft guidelines (e.g., coding styles).
-Patterned code generation can provide structured alternatives, such as replacing conditional statements with lookup tables or state machines.
-Integrating a naming convention engine allows automated conversion of variable and class names (e.g., transforming snake_case to PascalCase).
-Additionally, leveraging Abstract Syntax Tree (AST) analysis helps detect and transform prohibited structures, such as replacing for loops with while loops or recursive calls.
-Finally, conflict resolution mechanisms can be introduced to handle contradictory user instructions, such as providing alternative solutions when certain language features are unavailable (e.g., requesting switch-case in Python).
+Analysis of model-generated responses shows frequent deviations from instructions, especially in naming conventions and formatting constraints.
+Models often ignore global formatting rules, such as line limits, and inconsistently follow naming conventions.
+For example, when instructed to use PascalCase, models sometimes output lowercase or underscore-separated formats (e.g., incorrectly transforming current_power into CurrentPower).
+A notable issue is the disregard for prohibitive instructions.
+For instance, models often use if statements despite being instructed to avoid them in favor of ternary operators or data structures like dictionaries, revealing gaps in constraint enforcement.
+5.4 Improving Instruction Compliance
+Appendix Table 5 highlights strategies to improve adherence.
+Supervised Fine-Tuning (SFT) proves effective, especially in the Llama series, while larger models like Qwen2.5-72B-Instruct outperform smaller ones in instruction-following accuracy.
+Key improvements include prioritizing hard constraints (e.g., syntax rules) over soft guidelines (e.g., coding styles).
+Patterned code generation can replace conditional statements with lookup tables or state machines.
+A naming convention engine can automate variable name formatting (e.g., converting snake_case to PascalCase).
+Abstract Syntax Tree (AST) analysis can detect and transform prohibited structures, such as replacing for loops with while loops.
+Conflict resolution mechanisms can address contradictory instructions, offering alternative solutions when certain language features are unavailable (e.g., using Python’s alternatives to switch-case).
 6 Related Work
-Code generation and instruction-following are pivotal capacities under examination in AI research (Feng et al., 2020; Sun et al., 2024; Luo et al., 2024; Wang et al., 2023b; Kim et al., 2018; Li et al., 2023; Lu et al., 2021; Li et al., 2022; Wei et al., 2023; Nijkamp et al., 2023b; Zhuo et al., 2024; Jain et al., 2024; Nijkamp et al., 2023a; Zhang et al., 2023; Allal et al., 2023; Lozhkov et al., 2024a; Roziere et al., 2023; Lozhkov et al., 2024b; Wang et al., 2021; Yan et al., 2023).
+Code generation and instruction-following are pivotal capacities under examination in AI research Feng et al.
+(2020); Sun et al.
+(2024); Luo et al.
+(2024); Wang et al.
+(2023b); Kim et al.
+(2018); Li et al.
+(2023); Lu et al.
+(2021); Li et al.
+(2022); Wei et al.
+(2023); Nijkamp et al.
+(2023b); Zhuo et al.
+(2024); Jain et al.
+(2024); Nijkamp et al.
+(2023a); Zhang et al.
+(2023); Allal et al.
+(2023); Lozhkov et al.
+(2024a); Roziere et al.
+(2023); Lozhkov et al.
+(2024b); Wang et al.
+(2021); Yan et al.
+(2023).
 Several benchmarks have been devised to appraise these capabilities in large-scale models.
-For code generation, benchmarks like McEval (Chai et al., 2024), FullStackBench (Liu et al., 2024), Repocoder (Zhang et al., 2023), Repobench (Liu et al., 2023), and LiveCodeBench (Jain et al., 2024) have been notable.
-Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench (Qin et al., 2024), CFBench (Zhang et al., 2024), Instruct-following (Zhou et al., 2023), and FollowBench (Jiang et al., 2024), each tailored to assess different aspects of following instructions given to models.
+For code generation, benchmarks like McEval Chai et al.
+(2024), FullStackBench Liu et al.
+(2024), Repocoder Zhang et al.
+(2023), Repobench Liu et al.
+(2023), and LiveCodeBench Jain et al.
+(2024) have been notable.
+Similarly, instruction-following capacities are gauged through benchmarks such as InfoBench Qin et al.
+(2024), CFBench Zhang et al.
+(2024), Instruct-following Zhou et al.
+(2023), and FollowBench Jiang et al.
+(2024), each tailored to assess different aspects of following instructions given to models.
 7 Conclusion
@@
 It introduces novel evaluation metrics—Completely Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), Rigorous Satisfaction Rate (RSR), and Consistent Continuity Satisfaction Rate (CCSR)—to assess multi-constraint handling across multiple dimensions.
+8 Limitations
+Limited Language Support.
+CodeIF includes key languages like Java, Python, Go, and C++, but excludes popular ones like JavaScript, Ruby, and Swift.
+Expanding language coverage would improve its applicability in diverse contexts.
+Static Evaluation Focus.
+CodeIF focuses mainly on static code properties, such as structure and naming conventions, while overlooking dynamic factors like runtime behavior, performance, and debugging.
+Including dynamic evaluation would better reflect real-world development challenges.
+Uniform Metric Weighting.
+The metrics (CSR, SSR, RSR, CCSR) treat all constraints equally, which may not align with practical priorities.
+For example, syntactic correctness is often more critical than naming conventions.
+Introducing weighted scoring could enhance the interpretability of model performance.
+Appendix A Prompt Template
+Prompt for Instruction Generation You are an instruction compliance evaluator, required to assess the instruction compliance ability of large models.
+Therefore, you need to generate a series of data for the code generation instruction detection of large models.
+[Input Format]
+I will input a series of data, and you need to generate a dictionary based on these data, which includes two fields “question” and “instruction_list”
+Original question: {Original question} Original instruction list:
+{instruction_list}
+Input Explanation
+The original question is the original question. It contains the original code generation problem.
+The original instruction list is the original instruction list.
+It contains randomly generated code compliance instructions.
+Some instructions will contain directive keywords that need to be replaced and are wrapped in {{}}.
+Return Format
+Return a json data, do not have extra output. The returned dictionary contains two fields: “question” and “instruction_list”
+The format is as follows:
+    "question": "Optimized question",
+    "instruction_list": [
+            "instruction_id": "id1",
+            "instruction": "Instruction 1"
+Explanation
+“question”: It is the optimized question, which does not contain any directive instructions, only contains the explanation of the original question.
+It does not contain any restrictions on the code.
+Move the instructions in the question to the instruction list
+“instruction_list”: It is the optimized instruction list.
+You should optimize according to the meaning of the question.
+More in line with the meaning of the question.
+Instead of directly outputting the original instruction list, note that you should replace all directive keywords that need to be replaced and are wrapped in , and the final output should not contain directive keywords that need to be replaced.
+Generation Requirements
+question: Please generate the optimized question based on the following data, which does not contain any directive instructions, only contains the core content of the original question.
+instruction_list: Originated from the input original instruction list.
+If there are instructions that completely conflict with the meaning of the question or instructions that conflict with each other You should delete as little as possible, you should modify more.
+Please replace according to the content in the original instruction_list, you should delete as little as possible.
+Unless it is contradictory instructions, or instructions that cannot be achieved at all, if you only need to generate additional code to meet the requirements, you can replace it.
+Prompt for Code Generation As a programming assistant, your task is to generate code snippets based on the user question and instructions given below: Please consider the following points while generating the code snippet: - Make sure you follow the user instructions word to word.
+If the instruction says to use a specific language or a specific method, use exactly that.
+- Your output should be a valid code snippet in the programming language indicated in the question or the instructions.
+- Pay close attention to the syntax and formatting rules of the programming language that you are using.
+The code should be well-formatted and easy to read.
+- Make sure that the code snippet you are generating is efficient and is not overly complicated.
+Output Format: The output should be a valid, well-formatted, and efficient code snippet that adheres to the above question and instructions.
+Task information User Question: {question} Instructions: {instructions_str} Please generate the code snippet based on the above information:
+Prompt for Answer Judgment As a programming assistant, your task is to evaluate whether the generated code strictly follows the instructions given in light of the user’s problem and directives.
+You need to return a list of the same length as the instructions, containing only ’Yes’ and ’No’, indicating whether the model adhered to each specific instruction.
+Consider the following when making judgments: - You must strictly follow the user’s instructions.
+If the instruction requires the use of a specific language or method, you must explicitly check if the code utilizes it.
+- Your output should be a list of the same length as the instructions, containing only ’Yes’ and ’No’.
+- Pay close attention to the programming language syntax and formatting rules you are evaluating.
+The code should be neatly organized and easy to read.
+- The list you generate should be valid and not overly complex.
+Task Information User question:
+{question}
+Instructions:
+{instructions_str}
+Model-generated response:
+{generated_code}
+Based on the information provided, determine whether the model has followed the instructions, and return a list of the same length as the instructions, containing only ‘Yes’ and ‘No’.
+Please note!!!
+Your output should only contain the list, with no other content.
+The items in the list should only be ‘Yes’ and ‘No’, with no other words included.
+Appendix B More Resluts
+ID Type Instruction Format Format Keys
+1 global Your entire response should be written in {programming_language}, the use of other programming languages is not allowed. ["programming_language"]
+2 global Your code lines should not exceed {characters_num} characters. ["characters_num"]
+3 global Your code should use global variables. []
+4 global Your code should not use global variables. []
+5 global Your function should have at most {parameter_count} parameters. ["parameter_count"]
+6 global Your code should not have more than {function_count} functions. ["function_count"]
+7 global Your code should not have more than {class_count} classes. ["class_count"]
+8 global Your code should not use the {keyword} keyword. ["keyword"]
+9 global Your function should not exceed {line_num} lines. ["line_num"]
+10 global Your answer in total should not exceed {line_num} lines. ["line_num"]
+11 global Your code should use the {keyword} keyword. ["keyword"]
+12 structural control Your code should use data structure {data_structure}. ["data_structure"]
+13 structural control Your code should not use data structure {data_structure}. ["data_structure"]
+14 structural control Your code should use for-loop. []
+15 structural control Your code should not use for-loop. []
+16 structural control Your code should use while-loop. []
+17 structural control Your code should not use while-loop. []
+18 structural control Your code should use if statement for decision making. []
+19 structural control Your code should not use if statement for decision making. []
+20 structural control Your code should use switch statement for decision making. []
+21 structural control Your code should not use switch statement for decision making. []
+22 variable Your code should define a variable named {variable_name}. ["variable_name"]
+23 variable Your code should define an enumeration named {enumeration_name} ["enumeration_name"]
+24 variable The variable names in your code should follow the {naming_convention} naming convention ["naming_convention"]
+25 variable Variable {variable_name}, type should be {variable_type}. ["variable_name", "variable_type"]
+26 variable Variable {variable_name}, should be a global variable. ["variable_name"]
+27 variable Variable {variable_name}, should not be a global variable. ["variable_name"]
+28 variable Variable {variable_name}, the initial value should be {variable_value}. ["variable_name", "variable_value"]
+29 variable Variable {variable_name}, should be a constant. ["variable_name"]
+30 variable Variable {variable_name} should not be a constant. ["variable_name"]
+31 function Your code should include a function named {function_name}. ["function_name"]
+32 function The function names in your code should follow the {naming_convention}. naming convention ["naming_convention"]
+33 function Your code should not use any functions from the {disallowed_function_list}. ["disallowed_function_list"]
+34 interface Your code should define an interface named {interface_name}. ["interface_name"]
+35 interface The interface names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
+36 class Your code should define a class named {class_name}. ["class_name"]
+37 class The class names in your code should follow the {naming_convention} naming convention. ["naming_convention"]
+38 file Your code should be organized in a package named {package_name}. ["package_name"]
+39 file Your code should import the following libraries {library_list}. ["library_list"]
+40 file Your code should use the function {function_name} from the library {library_name}. ["function_name", "library_name"]
+41 file Your code should not use the following libraries {disallowed_library_list}. ["disallowed_library_list"]
+42 combination You should initialize an object named {object_name} as an instance of the {class_name} class using {parameters_name_list} for initialization.
+["object_name", "class_name", "parameters_name_list"]
+43 combination You should define an interface named {interface_name} that includes these methods {method_name_list}. ["interface_name", "method_name_list"]
+44 combination Your code should define a class named {class_name} that implements the {interface_name} interface. ["class_name", "interface_name"]
+45 combination In your code, the class {class_name} should have these properties {properties_name_list}. ["class_name", "properties_name_list"]
+46 combination In your code, the class {class_name} should have these methods {method_name_list}. ["class_name", "method_name_list"]
+47 combination The function {function_name} should take {parameter_name_list} as parameters. ["function_name", "parameter_name_list"]
+48 combination The function {function_name} should return a {return_type} as its result. ["function_name", "return_type"]
+49 combination Your code should be organized in a package named {package_name}, which should contain these classes {class_name_list}. ["package_name", "class_name_list"]
+50 combination Your code should be organized in a package named {package_name}, which should contain these functions {function_name_list}. ["package_name", "function_name_list"]
+Table 4: Constraint Instruction Table
+Models CSR SSR RSR CCSR
+Full Easy Hard Full Easy Hard Full Easy Hard Full Easy Hard
+Llama-3.2-1b-instruct 0.034 0.046 0.022 0.218 0.277 0.159 0.182 0.231 0.132 0.152 0.197 0.107
+Qwen2.5-1.5b-instruct 0.034 0.053 0.015 0.265 0.334 0.197 0.222 0.282 0.162 0.181 0.234 0.128
+Qwen2.5-coder-1.5b-instruct 0.058 0.086 0.03 0.358 0.436 0.281 0.301 0.371 0.233 0.251 0.314 0.189
+Qwen2.5-3b-instruct 0.078 0.109 0.046 0.415 0.489 0.34 0.357 0.432 0.282 0.299 0.364 0.233
+Llama-3.2-3b-instruct 0.101 0.137 0.065 0.396 0.473 0.318 0.344 0.419 0.268 0.305 0.375 0.235
+GPT-3.5-turbo 0.102 0.14 0.065 0.41 0.467 0.353 0.362 0.42 0.303 0.314 0.369 0.259
+Qwen2.5-coder-3b-instruct 0.097 0.142 0.051 0.445 0.529 0.359 0.383 0.464 0.301 0.33 0.401 0.258
+Llama-3.1-8b 0.129 0.178 0.08 0.452 0.551 0.353 0.402 0.497 0.306 0.352 0.44 0.263
+Llama-3.1-8b-instruct 0.145 0.187 0.102 0.467 0.544 0.388 0.418 0.493 0.34 0.37 0.444 0.295
+Qwen2.5-coder-7b-instruct 0.142 0.198 0.087 0.514 0.59 0.438 0.453 0.533 0.373 0.39 0.463 0.318
+Ministral-3b 0.127 0.162 0.092 0.526 0.591 0.46 0.458 0.527 0.39 0.4 0.458 0.342
+Phi-3.5-mini-128k-instruct 0.154 0.217 0.09 0.514 0.635 0.391 0.456 0.574 0.337 0.405 0.51 0.299
+Qwen2.5-7b-instruct 0.153 0.201 0.104 0.535 0.599 0.471 0.475 0.546 0.405 0.416 0.479 0.353
+Ministral-8b 0.161 0.205 0.116 0.552 0.614 0.489 0.486 0.552 0.419 0.431 0.49 0.371
+Gemma-2-9b-it 0.171 0.21 0.131 0.573 0.642 0.504 0.513 0.587 0.44 0.445 0.508 0.383
+Llama-3.1-70b 0.196 0.232 0.16 0.61 0.664 0.555 0.545 0.607 0.482 0.482 0.533 0.43
+Qwen2.5-coder-14b-instruct 0.218 0.276 0.16 0.596 0.667 0.525 0.539 0.614 0.463 0.483 0.55 0.416
+Qwen2.5-14b-instruct 0.238 0.279 0.198 0.61 0.676 0.543 0.557 0.628 0.486 0.498 0.565 0.431
+Gemini-2.0-flash-exp 0.254 0.29 0.218 0.615 0.648 0.583 0.556 0.593 0.518 0.514 0.547 0.481
+Gemma-2-27b-it 0.245 0.3 0.19 0.658 0.709 0.607 0.596 0.652 0.54 0.533 0.588 0.478
+Llama-3.1-70b-instruct 0.265 0.3 0.229 0.675 0.723 0.627 0.612 0.667 0.556 0.559 0.601 0.516
+Qwen2.5-32b-instruct 0.294 0.337 0.251 0.68 0.722 0.638 0.621 0.674 0.568 0.56 0.604 0.515
+Qwen2.5-72b-instruct 0.281 0.319 0.244 0.685 0.734 0.634 0.621 0.677 0.564 0.569 0.619 0.518
+Codestral-2501 0.28 0.339 0.219 0.683 0.748 0.617 0.621 0.691 0.551 0.571 0.633 0.507
+Phi-4 0.312 0.361 0.262 0.698 0.735 0.66 0.635 0.681 0.589 0.589 0.631 0.546
+Llama-3.3-70b-instruct 0.307 0.359 0.255 0.698 0.749 0.647 0.632 0.691 0.574 0.589 0.643 0.536
+GPT-4o-mini 0.292 0.348 0.237 0.731 0.78 0.682 0.665 0.724 0.606 0.609 0.66 0.559
+GPT-4o 0.338 0.392 0.283 0.721 0.77 0.671 0.665 0.721 0.609 0.616 0.668 0.563
+Qwen2.5-coder-32b-instruct 0.365 0.422 0.307 0.736 0.767 0.704 0.679 0.723 0.635 0.634 0.669 0.599
+Gemini-exp-1206 0.357 0.41 0.303 0.744 0.781 0.707 0.685 0.734 0.636 0.636 0.675 0.597
+Gemini-1.5-pro 0.351 0.383 0.318 0.763 0.794 0.732 0.704 0.744 0.663 0.647 0.679 0.615
+GPT-4o-2024-11-20 0.383 0.441 0.325 0.748 0.792 0.702 0.689 0.745 0.633 0.65 0.698 0.602
+Claude-3-5-sonnet-20241022 0.444 0.525 0.362 0.727 0.784 0.669 0.692 0.757 0.626 0.652 0.715 0.587
+Deepseek-coder 0.41 0.45 0.37 0.805 0.836 0.773 0.749 0.791 0.707 0.699 0.732 0.666
+Deepseek-v3 0.414 0.468 0.359 0.821 0.847 0.794 0.764 0.806 0.723 0.712 0.743 0.68
+Table 5: CodeIF evaluation results of different difficulties. We use bold font to mark the best results in all models.
+Models Global Structural Control Variable Interface Function Class File Combination
+Llama-3.2-1b-instruct 0.186 0.190 0.206 0.144 0.284 0.260 0.198 0.172
+Qwen2.5-1.5b-instruct 0.244 0.236 0.221 0.213 0.355 0.315 0.230 0.213
+Qwen2.5-coder-1.5b-instruct 0.328 0.304 0.326 0.293 0.436 0.426 0.351 0.304
+Qwen2.5-3b-instruct 0.383 0.346 0.412 0.383 0.468 0.481 0.383 0.366
+Llama-3.2-3b-instruct 0.344 0.332 0.393 0.376 0.454 0.447 0.363 0.367
+GPT-3.5-turbo 0.388 0.344 0.417 0.375 0.467 0.449 0.378 0.352
+Qwen2.5-coder-3b-instruct 0.397 0.367 0.438 0.419 0.511 0.507 0.415 0.403
+Llama-3.1-8b 0.410 0.355 0.451 0.424 0.500 0.503 0.413 0.413
+Llama-3.1-8b-instruct 0.422 0.373 0.482 0.455 0.524 0.499 0.407 0.437
+Qwen2.5-coder-7b-instruct 0.479 0.419 0.497 0.502 0.576 0.571 0.492 0.487
+Ministral-3b 0.472 0.403 0.527 0.512 0.618 0.609 0.524 0.535
+Phi-3.5-mini-128k-instruct 0.461 0.410 0.512 0.531 0.562 0.574 0.485 0.491
+Qwen2.5-7b-instruct 0.484 0.425 0.532 0.548 0.616 0.591 0.520 0.520
+Ministral-8b 0.497 0.433 0.541 0.570 0.622 0.631 0.527 0.557
+Gemma-2-9b-it 0.541 0.498 0.599 0.510 0.659 0.618 0.543 0.511
+Llama-3.1-70b 0.558 0.500 0.652 0.653 0.685 0.671 0.545 0.597
+Qwen2.5-coder-14b-instruct 0.541 0.467 0.623 0.669 0.645 0.652 0.547 0.594
+Qwen2.5-14b-instruct 0.569 0.526 0.652 0.592 0.649 0.644 0.533 0.559
+Gemini-2.0-flash-exp 0.555 0.526 0.653 0.666 0.685 0.669 0.564 0.615
+Gemma-2-27b-it 0.621 0.569 0.699 0.640 0.722 0.710 0.607 0.637
+Llama-3.1-70b-instruct 0.606 0.546 0.722 0.718 0.744 0.738 0.603 0.680
+Qwen2.5-32b-instruct 0.637 0.581 0.713 0.712 0.732 0.742 0.601 0.653
+Qwen2.5-72b-instruct 0.633 0.570 0.734 0.711 0.727 0.726 0.645 0.686
+Codestral-2501 0.617 0.552 0.723 0.718 0.733 0.746 0.651 0.694
+Phi-4 0.633 0.586 0.734 0.739 0.721 0.752 0.677 0.710
+Llama-3.3-70b-instruct 0.621 0.634 0.733 0.730 0.759 0.738 0.645 0.695
+GPT-4o-mini 0.671 0.663 0.787 0.774 0.784 0.783 0.657 0.710
+GPT-4o 0.665 0.651 0.742 0.759 0.743 0.759 0.666 0.716
+Qwen2.5-coder-32b-instruct 0.683 0.654 0.776 0.763 0.772 0.758 0.695 0.736
+Gemini-exp-1206 0.690 0.677 0.780 0.789 0.798 0.809 0.675 0.727
+Gemini-1.5-pro 0.718 0.696 0.814 0.800 0.812 0.815 0.672 0.749
+GPT-4o-2024-11-20 0.685 0.666 0.784 0.786 0.779 0.785 0.706 0.755
+Claude-3-5-sonnet-20241022 0.677 0.678 0.750 0.736 0.742 0.730 0.640 0.692
+Deepseek-coder 0.759 0.714 0.850 0.856 0.846 0.847 0.754 0.813
+Deepseek-v3 0.780 0.732 0.866 0.876 0.866 0.873 0.762 0.831
+Table 6: The performance of various models on CodeIF for different types of instructions
+Models CPP Java Python Go
+CCS CS SS RS CCS CS SS RS CCS CS SS RS CCS CS SS RS
+Llama-3.2-1b-instruct 0.123 0.023 0.185 0.150 0.190 0.037 0.265 0.221 0.179 0.047 0.262 0.223 0.086 0.022 0.117 0.096
+Qwen2.5-1.5b-instruct 0.171 0.023 0.250 0.206 0.191 0.026 0.277 0.228 0.197 0.047 0.298 0.257 0.151 0.040 0.216 0.179
+Qwen2.5-coder-1.5b-instruct 0.253 0.068 0.348 0.297 0.259 0.055 0.375 0.308 0.263 0.060 0.380 0.328 0.218 0.049 0.309 0.255
+Qwen2.5-3b-instruct 0.251 0.046 0.367 0.302 0.310 0.078 0.419 0.367 0.306 0.092 0.435 0.384 0.327 0.093 0.433 0.365
+Llama-3.2-3b-instruct 0.284 0.073 0.377 0.313 0.345 0.121 0.435 0.380 0.321 0.112 0.429 0.383 0.244 0.084 0.304 0.265
+GPT-3.5-turbo 0.301 0.085 0.388 0.332 0.367 0.134 0.461 0.409 0.265 0.092 0.371 0.334 0.318 0.088 0.412 0.363
+Qwen2.5-coder-3b-instruct 0.339 0.103 0.444 0.380 0.338 0.101 0.453 0.391 0.320 0.091 0.446 0.390 0.323 0.093 0.431 0.363
+Llama-3.1-8b 0.319 0.115 0.409 0.354 0.366 0.130 0.477 0.420 0.376 0.152 0.485 0.446 0.330 0.110 0.413 0.363
+Llama-3.1-8b-instruct 0.328 0.112 0.432 0.375 0.408 0.173 0.503 0.447 0.393 0.147 0.496 0.455 0.325 0.133 0.406 0.365
+Qwen2.5-coder-7b-instruct 0.389 0.147 0.505 0.434 0.375 0.118 0.503 0.444 0.400 0.155 0.531 0.475 0.401 0.154 0.516 0.456
+Ministral-3b 0.356 0.107 0.473 0.401 0.410 0.150 0.542 0.476 0.404 0.112 0.538 0.481 0.430 0.138 0.544 0.464
+Phi-3.5-mini-128k-instruct 0.354 0.108 0.461 0.388 0.426 0.179 0.532 0.478 0.440 0.180 0.559 0.510 0.380 0.131 0.482 0.422
+Qwen2.5-7b-instruct 0.401 0.162 0.514 0.448 0.439 0.152 0.559 0.495 0.397 0.147 0.523 0.471 0.429 0.154 0.541 0.485
+Ministral-8b 0.400 0.143 0.518 0.439 0.434 0.158 0.560 0.495 0.410 0.142 0.538 0.481 0.494 0.214 0.599 0.532
+Gemma-2-9b-it 0.446 0.200 0.560 0.499 0.446 0.164 0.576 0.518 0.380 0.131 0.510 0.456 0.542 0.204 0.678 0.609
+Llama-3.1-70b 0.487 0.201 0.598 0.518 0.507 0.232 0.632 0.572 0.425 0.136 0.579 0.521 0.522 0.226 0.635 0.571
+Qwen2.5-coder-14b-instruct 0.464 0.224 0.572 0.514 0.478 0.206 0.592 0.535 0.522 0.216 0.653 0.594 0.454 0.235 0.544 0.490
+Qwen2.5-14b-instruct 0.481 0.230 0.590 0.533 0.528 0.265 0.639 0.581 0.472 0.188 0.599 0.550 0.511 0.281 0.603 0.557
+Gemini-2.0-flash-exp 0.491 0.259 0.587 0.519 0.575 0.309 0.664 0.604 0.468 0.207 0.584 0.533 0.514 0.233 0.619 0.558
+Gemma-2-27b-it 0.529 0.271 0.645 0.579 0.551 0.261 0.676 0.616 0.465 0.179 0.604 0.543 0.611 0.289 0.727 0.665
+Llama-3.1-70b-instruct 0.535 0.267 0.653 0.578 0.581 0.276 0.685 0.620 0.555 0.251 0.696 0.639 0.557 0.264 0.655 0.596
+Qwen2.5-32b-instruct 0.551 0.314 0.655 0.602 0.589 0.330 0.706 0.638 0.522 0.231 0.665 0.609 0.580 0.311 0.690 0.634
+Qwen2.5-72b-instruct 0.543 0.297 0.638 0.574 0.580 0.288 0.701 0.633 0.574 0.284 0.702 0.651 0.573 0.249 0.687 0.610
+Codestral-2501 0.562 0.307 0.658 0.595 0.583 0.301 0.694 0.632 0.566 0.249 0.693 0.637 0.569 0.261 0.681 0.611
+Phi-4 0.570 0.331 0.663 0.601 0.612 0.328 0.719 0.650 0.587 0.295 0.714 0.660 0.577 0.292 0.679 0.613
+Llama-3.3-70b-instruct 0.558 0.300 0.652 0.582 0.621 0.348 0.713 0.644 0.572 0.264 0.709 0.653 0.602 0.317 0.712 0.640
+GPT-4o-mini 0.582 0.292 0.684 0.615 0.620 0.299 0.738 0.667 0.586 0.261 0.731 0.674 0.661 0.330 0.775 0.707
+GPT-4o 0.600 0.337 0.698 0.639 0.652 0.368 0.748 0.693 0.600 0.312 0.723 0.676 0.603 0.332 0.701 0.636
+Qwen2.5-coder-32b-instruct 0.633 0.384 0.717 0.658 0.654 0.401 0.753 0.699 0.621 0.342 0.736 0.688 0.624 0.322 0.731 0.661
+Gemini-exp-1206 0.640 0.424 0.726 0.672 0.650 0.360 0.755 0.689 0.590 0.290 0.724 0.674 0.677 0.373 0.777 0.710
+Gemini-1.5-pro 0.635 0.370 0.741 0.676 0.674 0.379 0.783 0.720 0.610 0.278 0.758 0.706 0.674 0.395 0.764 0.707
+GPT-4o-2024-11-20 0.653 0.374 0.741 0.669 0.683 0.434 0.776 0.716 0.612 0.355 0.724 0.682 0.653 0.358 0.747 0.683
+Claude-3-5-sonnet-20241022 0.615 0.425 0.684 0.643 0.720 0.504 0.789 0.749 0.611 0.396 0.703 0.674 0.650 0.444 0.716 0.686
+Deepseek-coder 0.709 0.441 0.802 0.735 0.731 0.463 0.819 0.764 0.657 0.336 0.791 0.747 0.702 0.403 0.805 0.744
+Deepseek-v3 0.725 0.435 0.831 0.762 0.753 0.497 0.839 0.787 0.651 0.315 0.793 0.744 0.722 0.404 0.822 0.76
+Table 7: the evaluation results of different languages on CODEIF.
+The metrics include Consistent Continuity Satisfaction Rate (CCSR), Complete Satisfaction Rate (CSR), Soft Satisfaction Rate (SSR), and Rigorous Satisfaction Rate (RSR).
+Appendix C Baselines
+We evaluate over 30 language models spanning both open-source architectures and commercial APIs.
+The Meta Llama 3 Series Touvron et al.
+(2023) contains Llama-3.2-1B/3B/8B/70B-Instruct variants and Llama-3.3-70B-Instruct.
+Qwen2.5 Series Yang et al.
+(2024) encompasses Qwen2.5-1.5B/3B/7B/14B/32B/72B-Instruct with dedicated code generation models Qwen2.5-Coder-1.5B/3B/7B/14B/32B-Instruct Hui et al.
+(2024).
+Mistral Series Jiang et al.
+(2023) includes Mistral-3B, Mistral-8B, and the code-specialized Codestral-2501.
+The evaluation covers Microsoft’s Phi-3.5-Mini-128K-Instruct (3.8B) and Phi-4 Abdin et al.
+(2024), along with Google’s Gemma-2-9B/27B-It Team (2024b).
+DeepSeek Series incorporates DeepSeek-Coder Guo et al.
+(2024) and DeepSeek-V3 DeepSeek-AI (2024).
+Commercial APIs include OpenAI’s GPT-3.5-Turbo, GPT-4O-Mini, GPT-4O-2024-05-13, and GPT-4O-2024-11-20 Achiam et al.
+(2023); Google’s Gemini-2.0-Flash-Exp, Gemini-Exp-1206, and Gemini-1.5-Pro Team (2024a); plus Anthropic’s Claude-3.5-Sonnet-20241022.
+Appendix D More Data Analysis
+Figure 7: Distribution of atomic instruction list lengths across difficulty levels.
+Figure 7 shows the proportion of each instruction category.
+Global constraints dominate (22.77%), followed by Variable constraints (18.17%).
+This distribution reflects CodeIF’s balanced focus on high-level structural coherence and fine-grained variable precision, ensuring comprehensive evaluation of code generation capabilities.
+Figure 8 compares instruction distribution across difficulty levels.
+Figure 8: The distribution of constraint instruction list lengths in datasets of different difficulties.